OpenStudy (anonymous):

Find the coordinates of the midpoint of AB. A(12, -7) and B(-4, 7). A. (-8, 7) B. (4,0) C. (8, -7) D. (8, 0)

OpenStudy (anonymous):

i think you should use the mid point formula to get the answer.

OpenStudy (rmrjr22):

\[(x2+x1)/2+(y2+y1)/2\]

OpenStudy (rmrjr22):

+ is supposed to be a ,

OpenStudy (anonymous):

use the formula @rmrjr22 wrote and get the answer,,

OpenStudy (anonymous):

its distance formula?

OpenStudy (anonymous):

B

OpenStudy (anonymous):

4,0

OpenStudy (rmrjr22):

midpoint

OpenStudy (rmrjr22):

no

OpenStudy (anonymous):

no no.. not distance formula.. the mid point formula..... the x- coordinate of the mid point =\[\frac{ x1+x2 }{ 2 }\] x1 - 12 x2 = -4 now, can you solve for the x- coordinate of mid point?

OpenStudy (anonymous):

((12 - 4) / 2, (-7 + 7) /2) = (8/2, 0/2) = (4, 0)

OpenStudy (anonymous):

Yes

OpenStudy (rmrjr22):

good

OpenStudy (anonymous):

The answer is B (4,0)

OpenStudy (anonymous):

yes i got b too

OpenStudy (rmrjr22):

yes

OpenStudy (anonymous):

ok thank you